OpenStudy (anonymous):

@pooja195 @Leader @sleepyjess

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (anonymous):

a?

OpenStudy (sleepyjess):

I know that TSO went over this in one of the previous questions. We need the opposite slope. If we have -x at the beginning, the slope is -1 which will make the line parallel to the original line.

OpenStudy (anonymous):

ok @sleepyjess

OpenStudy (sleepyjess):

So instead we need x

OpenStudy (anonymous):

ok

OpenStudy (sleepyjess):

Then, we need to figure out the next points from the given points and slope. To do that, we add 1 to each x and y value until x = 0

OpenStudy (sleepyjess):

So we start with (-4, -2), add 1 to each coordinate, we get (-3, -1).

OpenStudy (anonymous):

ok @sleepyjess, one question a was wrong xd

OpenStudy (sleepyjess):

huh?

OpenStudy (sleepyjess):

So what is the correct answer for this one with the information I have given?

OpenStudy (anonymous):

c?

OpenStudy (sleepyjess):

C still has -x

OpenStudy (sleepyjess):

We need x

OpenStudy (anonymous):

either b or d then @sleepyjess

OpenStudy (sleepyjess):

Yes, and what is the y-intercept from adding 1 until x = 0?

OpenStudy (anonymous):

d? then

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

YEP

OpenStudy (anonymous):

The first five terms of a sequence are shown below: 4, 7, 10, 13, 16 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 3n - 1; n ≥ 1 f(n) = 3n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1

OpenStudy (anonymous):

b?

OpenStudy (sleepyjess):

Yes

OpenStudy (anonymous):

If f(x) = 2x2 + 2, find f(5). 102 52 22 -1.5

OpenStudy (anonymous):

b?

OpenStudy (sleepyjess):

2(5)^2 + 2 10^2 + 2

OpenStudy (anonymous):

102?

OpenStudy (sleepyjess):

yes

OpenStudy (anonymous):

are you sure

OpenStudy (sleepyjess):

yes

OpenStudy (anonymous):